How would you solve this problem? The perimeter of the cover of an address book is 14 inches, and the diagonal measures 5 inches. What are the length and width of the cover?
Answer by arunpaul(104)

When the diagonal is 5 inch then the other two sides must be 3inch and 4 inch
so that 3^2+4^2 = 5^2
so the length will be 4inch and width will be 3 inch
so the perimeter will be
2(3+4)=14 inch (Ans)